1.05 L of ethane gas (C2H6) at 20.5*C and 771 mmHg burns in a bomb calorimeter, which makes the temperature rise to 100.5*C. What is the heat capacity of the calorimeter in kJ/K?
1
Expert's answer
2019-11-11T07:49:39-0500

Heat capacity HC=heat energy supplied E /rise in temperature T.

"HC=E\/T"

The heat energy supplied to the calorimeter is given by; E="m*cp*delta (T)"

1.05L of ethane=1050g=1.05kg

Specific heat capacity of ethane, cp= 0.69933(T/K)

Delta(T)=100.5-20.5= 80K

Therefore, E= 1.05*0.69933*80= 58.74372 joules=0.05874372KJ

The heat capacity of calorimeter is given by;

HC=0.05874372/80 = 7.343*10^-4 KJ/K